OpenStudy (anonymous):

How to find the derivative of 1/sqrt (x+2) using the chain rule?

OpenStudy (mathmale):

You'll find this problem easier if you re-write it first. 1/sqrt (x+2) should be written as y = (x+2)^(-1/2). Can you explain where the (-1/2) came from?

OpenStudy (anonymous):

yes you just rewrote the square root and its a negative because the square root is on the bottom.

OpenStudy (mathmale):

Yes. Next preliminary question: what is the derivative of x+2?

OpenStudy (anonymous):

its 1

OpenStudy (mathmale):

That's right. Good. Now, yo u have y=(x+2)^(-1/2). Use the power rule (first) followed by the chain rule (second) to find the derivative. Have you done this before? "Power rule with chain rule."

OpenStudy (anonymous):

uhm yes. this is what I got. -1/2(x+2)^-3/2

OpenStudy (mathmale):

Looks really good, EXCEPT that y ou must enclose the initial -1/2 within parentheses. Try again, please. Label your result, remembering that it's a derivative.

OpenStudy (anonymous):

(-1/2)(x+2)^-3/2

OpenStudy (anonymous):

thanks. I just struggle with rewriting the problem i guess. but thanks for your help!

OpenStudy (mathmale):

You're very welcome. Good luck to you!